The sum of linda's and steve's ages is 113. Steve is 11 years older than linda. How old is linda and how old is steve?
Answer by mananth(16946)

L+S =113
L=113-S
..
S=L+11
plug value of L
S=113-S+11
2S=124
S=62 ... Steve's age
Linda's age = 113-62=51 years